OpenStudy (anonymous):

Cards numbered 1, 1, 2, 3, 4, 4, 4, 5, 6, 6, 7, 8, 8, 8, and 9 are placed in a bag. If a card is drawn from the bag and then replaced 135 times, about how many times would a 6 be drawn? A. 2 B. 9 C. 15 D. 18

OpenStudy (kendricklamar2014):

How many 6's are there @shelbydancer

OpenStudy (anonymous):

2

OpenStudy (kendricklamar2014):

Yes, now, How many Total cards are there??

OpenStudy (kendricklamar2014):

@shelbydancer

OpenStudy (kendricklamar2014):

There are a total of 15 cards

OpenStudy (kendricklamar2014):

The probability of choosing a 6 would be:

OpenStudy (kendricklamar2014):

Each cards is replaced 135 so: 2/15 * 135 = 270/15 =

OpenStudy (kendricklamar2014):